Ms. Maili Sirrus obtained a balloon loan of $238,665 at a nominal rate of interest of 4.15 percent for 4 years. If repayment is scheduled at the end of every fifth day, obtain the level of the first few payments, assuming he will pay $ 36,223 at the end of the term.

Round your final answer to 2 decimal places.
Answer #1

Loan amount = 238,665; balloon payment = 36,223

Number of payments = number of years*number of days in a year/5 = (4*365)/5 = 292

Compounding rate = annual rate/number of payments in a year = 4.15%/(365/5) = 0.0568%

PV = -238,665; N = 292; rate = 0.0568%, solve for PMT.

Level payment every 5th day = $887.29